OpenStudy (anonymous):

At breakfast, Gino drank 2/3 of the orange juice in a container. He drank 15 ounces of orange juice. A. Write a multiplication equation that you can use to find the amount of orange juice that was in the container before Gino had breakfast. B. Solve the equation in Part A. Show your work.

OpenStudy (anonymous):

@icecreamman

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

OJ(2/3) = 15 OJ = 15 / (2/3) OJ = 15 * 3/2 OJ = ?

OpenStudy (anonymous):

27.5 oz i think

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

did that help? like did you get what i did?

OpenStudy (anonymous):

C. Tell how much orange juice was in the container before Gino had breakfast.

OpenStudy (anonymous):

can i get a medal first?

OpenStudy (vera_ewing):

What do you not understand about the question?